I would appreciate help in the following problem: w = P +1 over 2, for 1. Solve.
Thanking you in advance.

Answer by jb007mi6(53)

w=(P+1)/2
multiply both sides by 2 to clear fraction
2w=P+1
subtract P from both sides
2w-P=1
turn around for answer
1=2W-P